Please solve, show work, and give detail explanation6. A firm offers a 6.4 percent bond with a current market price of $760.50. The yield-to- maturity (YTM) is 8.52 percent. The
Answer #1
6) Years to maturity = =nper(rate,pmt,-pv,fv)/2
=               39 Years
Where,
rate = 8.52%/2 = 0.0426
pmt = 1000*6.4%*6/12 = 32
pv = $     760.50
fv = $ 1,000.00
7)
As per fisher equation,
(1+R) = (1+r)*(1+i) Where,
(1+0.075) = (1+r)*(1+0.02) R = Nominal rate =         0.075
1.075 = (1+r)*1.02 r = real rate = ?
1.053922 = 1+r i = Inflation rate =           0.02
0.053922 = r
So, real rate of return is 5.39%